OpenStudy (lidia.12):

How do you solve Log3(three is little)(x+6)-Log3(three is little)x=2

OpenStudy (mrnood):

First: Use the rules of logs to simplify the left side \[\log _{3}x -\log _{3} y = \log _{3} \frac{ x }{y }\] write the equation using this rule

OpenStudy (mrnood):

THEN you need to solve this \[\log _{3}x = 2\] find x